OpenStudy (nuccioreggie):

Multiply two and four-fifths negative five and two-thirds . negative ten and eight-fifteenths negative fifteen and thirteen-fifteenths ten and eight-fifteenths fifteen and thirteen-fifteenths

OpenStudy (nuccioreggie):

@TuringTest

OpenStudy (nuccioreggie):

@misty1212

OpenStudy (misty1212):

\[\huge 2\tfrac{4}{5}\times \left(-5\tfrac{2}{3}\right)\]

OpenStudy (nuccioreggie):

yea

OpenStudy (misty1212):

you know how to turn these in to improper fractions?

OpenStudy (nuccioreggie):

no but is trhe answer c or b

OpenStudy (misty1212):

\[\frac{14}{5}\times \left(-\frac{17}{3}\right)\]

OpenStudy (nuccioreggie):

thats not an answer choice

OpenStudy (misty1212):

you gotta multiply silly

OpenStudy (misty1212):

\[-\frac{238}{15}\]

OpenStudy (nuccioreggie):

ok

OpenStudy (danjs):

ahah

OpenStudy (misty1212):

then write as a mixed number

OpenStudy (danjs):

repost the same question hoping someone will hurry to the answer

OpenStudy (danjs):

i typed the answer in the last post you did fyi

OpenStudy (misty1212):

eventually you will get \[-15\tfrac{13}{15}\]

OpenStudy (nuccioreggie):

that was wrong @misty1212